# Local Loss Optimization in the Infinite Width: Stable Parameterization of Predictive Coding Networks and Target Propagation

### v1

@@ A.2 Predictive Coding with single-shot SI
Additionally, it admits the NTK parametrization at
@@
for $l<L$ where $e_{L,{0}}=y-W_{L}v_{L-1,0}=:\delta_{L}$. To keep the notation concise, we set $M_{L}=1$ in this proof. A generalization for $M_{L}=\Theta(1)$ is possible. Next, we define
$\delta_{l<L}:=\partial f/\partial u_{l}.$ (S.12)
@@
In PC, in addition to the usual learning rate $\eta$, there also exists $\gamma$. Therefore, in addition to the infinitesimal update of the learning rate $\eta$ for the weight update, we also consider the infinitesimal inference step size $\gamma_{L}$. By applying the perturbation of $\gamma_{L}$ to Conditions A.1 and S.2, we derive
@@
It is noteworthy that we retain the zero-th order terms, namely, $\Delta U_{l}$ and $\Delta V_{l}$ in the conditions. This is because, even without the inference phase, parameter updates can progress while the internal states remain at their initialization. Therefore, even if the maximalization of the order is less than $\Theta(1)$ in the first-order perturbation terms, stable learning can still occur. Since $\mu$P aims to maximize the order of updates as much as possible, we require the first-order terms of Eqs. (S.21,S.22) to be $\Theta(1)$ whenever possible.
@@
in the infinite-width limit [Yang, 2020].
@@
where we used $v_{L-2,1}-v_{L-2,0}=-\gamma_{L-1}\gamma_{L}\delta_{L-2}$ from Eq. (S.13). Let us recall that a variable without an index indicates the initial state at $s=0$. The variable $\delta_{L-1}$ includes $W_{L}$ whereas $h_{L-1}$ is independent of it. Therefore, by applying the CLT with respect to $W_{L}$, we have
@@
Comparing the zero-th and first order terms (S.47,S.48), we obtain
@@
We can also derive the NTK parameterization, which is a commonly used term for the kernel regime for $r_{l<L}=1/2$ [Yang and Hu, 2021]:
@@
It is noteworthy that the gradient computed by Eq. (S.15) differs from $\delta_{l}$ in standard GD, implying that the NTK matrix also deviates from $\nabla_{\theta}f^{\top}\nabla_{\theta}f$. Even in this case, the NTK regime can emerge with a certain modified kernel composed of $e_{l}$ and $h_{l}$. A similar situation arises in the NTK regime of second-order optimization [Karakida and Osawa, 2020]. Altough the preconditioner modifies the NTK matrix, the linearization of the model still holds, allowing the emargence of the kernel regime.
A.3 Fixed points of PC in Linear Networks
A.3.1 Proof for Theorem 4.2
@@ A.3 Fixed points of PC in Linear Networks
Even for linear networks, the properties of the fixed points have rarely been analyzed. An exception is a recent study by Innocenti et al. [2024]. They explicitly derived the free energy at a fixed point to analyze the parameter loss landscape of a naive PC. However, their analysis uses an unfolding calculation of a hierarchical Gaussian model to directly derive the free energy. Although this is an elegant derivation, it is not a method for explicitly obtaining the fixed points themself. Additionally, since their proof is based on $\gamma=1$, we need another method to determine the dependence on the inference size. Here, we provide a derivation of the states at the fixed point that can be used more generally for various inference sizes and adding a nudge term (in Section A.3.3).
@@
Proof. One can easily derive this inverse matrix. A simple derivation is achived by induction. We can express
@@
By using Lemma A.6, we can transform Eq. (S.67) as follows:
@@
Altough this equation can not be solved explicitly for $v_{l-1}^{*}$, we can, nonetheless, solve it by summing over $e_{l}^{*}$ as follows:
@@
A.3.2 Balance condition determining $\gamma_{l<L}$
@@
A.3.3 Nudged Predictive Coding
Predictive Coding with a nudge term updates the state $v_{l}$ and weights $W_{l}$ to minimize the following free energy function [Alonso et al., 2022, Millidge et al., 2023, Pinchetti et al., 2024]:
Here, $\beta$ is a nudge coefficient parameter that covers some variants of the PC algorithms in the previous work. We can extend Theorem 4.2 to the nudged PC.
Theorem A.7.
@@ A.3.3 Nudged Predictive Coding
From Lemma A.6, the above equation is transformed into
@@
A.4 Target Propagation
A.4.1 Derivation of Theorem 19.
@@ A.4 Target Propagation
where $h_{l}=\phi(W_{l}h_{l-1})$. The feedback network is given by the network with Eq. (S.103). As a side note, this weight is essentially the same as the pseudo-inverse weight, which is known as an extension of the Hebbian weight [Kanter and Sompolinsky, 1987].
Local targets of DTP. DTP is an improved methods of TP, where $\hat{h}_{L}$ is propagated as follows:
@@
$\hat{h}_{l}$ $=g_{l+1}\left(\hat{h}_{l+1}\right)-\left(g_{l}\left(h_{l+1}\right)-h_{l}\right)$ (S.106)
@@
$=h_{l}-\beta\prod_{i=l+1}^{L}Q_{l}\delta_{L}.$ (S.110)
Therefore, at the equilibrium point for $Q_{l}$, for $l\leq L-1$, we have
$\hat{h}_{l}-h_{l}$ $=-\beta\prod_{i=l+1}^{L}Q_{l}^{*}\delta_{L}$ (S.111)
@@
where $\delta_{L}=\partial\mathcal{L}/\partial f$. To avoid an uninteresting change of order, we introduce $\mu_{l}=\mu_{l}^{\prime}/M^{\bar{\mu}_{l}}$ and require that it have the same order as $K_{l}^{A}$. This is essentially equivalent to the valid condition argued in Ishikawa and Karakida [2024], which requires the damping term to have the same order as the preconditioner in the second-order optimization. We note that we can take the ridge-less limit $\mu_{l}^{\prime}\rightarrow 0+$ because $K^{A}_{l}$ (S.26) is typically set to be regular at random initialization in the neural tangent kernel literature [Jacot et al., 2018, Yang, 2020]. For instance, this holds true for normalized input samples with $\|x\|=1$.
@@
On Condition A.1.
@@
$=-\frac{1}{M^{\theta_{L-1}-1}}\beta h_{L-1}(K^{A}+\mu_{L}^{\prime}I)^{-1}(K^{A}-M^{-1}h_{L}^{\top}y)K^{A}_{L-1}.$ (S.119)
@@
$=-\frac{1}{M^{\theta_{l}-1}}\beta h_{l}\prod_{i=l+1}^{L}(h_{i}^{\top}h_{i}+\mu_{i}I)^{-1}h_{i}^{\top}h_{i}(h_{L}^{\top}h_{L}+\mu_{L}I)^{-1}h_{L}^{\top}\delta_{L}K^{A}_{l-1}$ (S.121)
$=-\frac{1}{M^{\theta_{l}-1}}\beta h_{l}\prod_{i=l+1}^{L}(K^{A}_{i}+\mu_{i}^{\prime}I)^{-1}K^{A}_{i}(K^{A}_{L}+\mu_{L}^{\prime}I)^{-1}(K^{A}_{L}-M^{-1}h_{L}^{\top}y)K^{A}_{l-1}$ (S.122)
@@
$\partial_{\eta^{\prime}}(W_{L,0}\Delta h_{L-1,1})\big{|}_{\eta^{\prime}=0}$ $=\frac{1}{M^{\theta_{L-1}}}W_{L,0}(\hat{h}_{L-1}-h_{L-1})h_{L-1}^{\top}$ (S.126)
$=-\frac{1}{M^{\theta_{L-1}-1}}\beta h_{L}(h_{L}^{\top}h_{L}+\mu_{L}^{\prime}I)^{-1}h_{L}^{\top}\delta_{L}K^{A}_{L-1}$ (S.127)
$=-\frac{1}{M^{\theta_{L-1}-1}}\beta h_{L}(K^{A}_{L}+\mu_{L}^{\prime}I)^{-1}(K^{A}_{L}-M^{-1}h_{L}^{\top}y)K^{A}_{L-1}.$ (S.128)
@@
A.4.2 Disappearance of Kernel regime
Corollary A.8.
@@ A.4.2 Disappearance of Kernel regime
Note that, precisely speaking, $r_{L-1}=0$ does not necessarily imply $r_{l<L-1}=0$. However, it is often considered unnatural (or uninteresting) to examine cases in which the progress of learning depends on individual layers. Therefore, the $\mu$P typically assumes a uniform parameterization, meaning $r_{l<L}=r$ (see Theorem G.4 of Yang and Hu [2021]). In this sense, $r_{L-1}=0$ indicates the disappearance of the kernel regime.
One might be surprised by the fact that $b_{L}=1/2$ is allowed in the feature learning and that the kernel regime disappears. Note that the feedback weight in the last layer (S.103) essentially differs from $W_{L}$. The feedback weight recieves $h_{L}$ as input whereas $W_{l}$ recieves $h_{L-1}$. This makes
@@
$W_{L}\sim 1/M^{a_{L}+b_{L}}$. The gradient is proportional to $Q_{L}$ in TP and $W_{L}$ in BP. The feedback weight contributes more significantly to TP’s gradient when $a_{L}+b_{L}\geq 1/2$. This eventually makes the index $r$ of the hidden layer (S.125) get quite large even for $a_{L}+b_{L}=1/2$. We also need to be careful about the order of condition A.2 (S.130). Because the feedback weight (S.103) has a lower alignment exponent [Everett et al., 2024], this causes the condition 2 of TP to be smaller than that of SGD (or K-FAC), i.e., $1/M^{r_{L-1}+(a_{L}+b_{L})-1}$. Therefore, the stable feature learning is possible even for $a_{L}+b_{L}=1/2$.
@@
Substituting these into Eqs. (S.119,S.122), we obtain
@@
A.5 Stable parametrization for feedback network
@@
where dividing by $M_{L-1}$ is to ensure that $L(Q_{l})=\Theta(1)$, which is the default setting in PyTorch. We consider the parametrization in the feedback network:
@@
Appendix B Additional Experiments
B.1 Predictive Coding
B.1.1 Linear Network
Figure S.1 : During training of the linear network, it converges to the analytical solution. We trained a 3-layer linear network using synthetic data.

### v2

@@ B.1 Derivation for Predictive Coding with single-shot SI
Additionally, it admits the NTK parameterization at
@@
for $l<L$ where $e_{L,{0}}=y-W_{L}v_{L-1,0}=:\delta_{L}$. When the CE loss is used instead of MSE loss, $\delta_{L}=y-f$ becomes $\delta_{L}=y-\text{softmax}(f)$, and the order analysis remains unchanged. To keep the notation concise, we set $M_{L}=1$ in this proof. A generalization for $M_{L}=\Theta(1)$ is possible. Next, we define
$\delta_{l<L}:=\partial u_{L}/\partial u_{l}.$ (S.25)
@@
In PC, in addition to the usual learning rate $\eta$, there also exists $\gamma$. Therefore, in addition to the infinitesimal update of the learning rate $\eta$ for the weight update, we also consider the infinitesimal inference step size $\gamma_{L}$. By applying the perturbation of $\gamma_{L}$ to Conditions A.3 and S.17, we derive
@@
It is noteworthy that we retain the zero-th order terms, namely, $\Delta U_{l}$ and $\Delta V_{l}$ in the conditions. This is because, even without the inference phase, parameter updates can progress while the internal states remain at their initialization. Therefore, even if the maximalization of the order is less than $\Theta(1)$ in the first-order perturbation terms, stable learning can still occur. Since $\mu$P aims to maximize the order of updates as much as possible, we require the first-order terms of Eqs. (S.34,S.35) to be $\Theta(1)$ whenever possible.
@@
in the infinite-width limit (Yang, 2020).
@@
where we used $v_{L-2,1}-v_{L-2,0}=-\gamma_{L-1}\gamma_{L}\delta_{L-2}$ from Eq. (S.26). Let us recall that a variable without an index indicates the initial state at $s=0$. The variable $\delta_{L-1}$ includes $W_{L}$ whereas $h_{L-1}$ is independent of it. Therefore, by applying the Central Limit Theorem with respect to $W_{L}$, we have
@@
Comparing the zero-th and first order terms (S.60,S.61), we obtain
@@
The above $\mu$P case assumes $a_{l}=0$. It is important to note that there is no issue in replacing $c_{l}$ with $\theta_{l}=2a_{l}+c_{l}$, which introduces an indeterminacy of $a_{l}=a_{l}+\alpha$ and $c_{l}=c_{l}-2\alpha$.
We can also derive the NTK parameterization, which is a commonly used term for the kernel regime for $r_{l<L}=1/2$ (Yang & Hu, 2021):
@@
It is noteworthy that the gradient computed by Eq. (S.28) differs from $\delta_{l}$ in standard SGD, implying that the NTK matrix also deviates from $\nabla_{\theta}f^{\top}\nabla_{\theta}f$. Even in this case, the NTK regime can emerge with a certain modified kernel composed of $e_{l}$ and $h_{l}$. A similar situation arises in the NTK regime of second-order optimization (Karakida & Osawa, 2020). Although the preconditioner modifies the NTK matrix, the linearization of the model still holds, allowing the emergence of the kernel regime.
B.2 Fixed points of PC in Linear Networks
B.2.1 Proof for Theorem 4.2
@@ B.2 Fixed points of PC in Linear Networks
Even for linear networks, the properties of the fixed points have rarely been analyzed. An exception is a recent study by Innocenti et al. (2024). They explicitly derived the free energy at a fixed point to analyze the parameter loss landscape of a naive PC. However, their analysis uses an unfolding calculation of a hierarchical Gaussian model to directly derive the free energy. Although this is an elegant derivation, it is not a method for explicitly obtaining the fixed points themself. Additionally, since their proof is based on $\gamma=1$, we need another method to determine the dependence on the inference size. Here, we provide a derivation of the states at the fixed point that can be used more generally for various inference sizes and add a nudge term (in Section B.2.4).
@@
Proof. One can easily derive this inverse matrix. A simple derivation is achieved by induction. We can express
@@
By using Lemma B.2, we can transform Eq. (S.80) as follows:
@@
Although this equation can not be solved explicitly for $v_{l-1}^{*}$, we can, nonetheless, solve it by summing over $e_{l}^{*}$ as follows:
@@
B.2.2 Proof for Corollary 4.3
Proof. We consider the order of $W_{L:i}W_{L:i}^{\top}$ in $C_{\gamma}(W)=\sum_{i=2}^{L}\frac{\gamma_{L}}{\gamma_{i-1}}W_{L:i}W_{L:i}^{\top}$. Note that computing the vector $v=W_{L:i}^{\top}$ is equivalent to signal propagation in a deep linear network with random weights $W_{l:L-1}^{T}$ and an input vector $W_{L}^{\top}$. Therefore, as in Eq. (S.39), we can evaluate $\|v\|_{2}^{2}$ using kernel computations in existing studies of random neural networks. Specifically, this is equivalent to computing $h_{l}^{\top}h_{l}$ in the following random neural network:
$u_{k}$ $=W_{k}h_{k-1},\quad h_{k}=\phi(u_{k})\quad(k=1,\dots,l),$ (S.96)
$x_{i}\sim\mathcal{N}(0,1/M^{a_{L}+b_{L}})\quad(i=1,...,M),$ (S.97)
$W_{k,ij}\sim\mathcal{N}(0,1/\sqrt{M})\quad(i,j=1,...,M).$ (S.98)
From Theorem 7.2 in Yang (2020), for any $x$, $q_{k}=h_{k}^{\top}h_{k}/M$ converges almost surely to
[…]
in the infinite-width limit. Since we are considering a linear activation function, we have $q_{l}=q_{l-1}$ and
$q_{l}=\frac{1}{M}x^{\top}x.$ (S.100)
Since $x$ is an i.i.d. Gaussian vector, the law of large numbers leads to $q_{l}$ of $\Theta(1/M^{2(a_{L}+b_{L})})$. For random weights given by $\mu$P i.e., $a_{L}+b_{L}=1$, we find $W_{L:i}W_{L:i}^{\top}=\Theta(1/M)$. Consequently, for ${\gamma}_{L}=\Theta(M)$, $C_{\gamma}(W)$ remains of order 1. In contrast, for $\gamma_{L}=\Theta(1)$, $C_{\gamma}(W)$ is $O(1/M)$, causing $I+C_{\gamma}(W)$ to approach $I$ in the infinite-width limit. As a result, $e_{l}^{*}=\delta_{l}$ and $v_{l}^{*}=u_{l}$, and the gradient in the infinite-width limit matches that of first-order gradient descent.
B.2.3 Balance condition determining $\gamma_{l<L}$
@@
B.2.4 Nudged Predictive Coding
We can extend Theorem 4.2 to the nudged PC.
Theorem B.3.
@@ B.2.4 Nudged Predictive Coding
From Lemma B.2, the above equation is transformed into
@@
Appendix C $\mu$P of Target Propagation
C.1 Derivation of Theorem 17
@@
where $h_{l}=\phi(W_{l}h_{l-1})$. The feedback network is given by the network with Eq. (S.120). As a side note, this weight is essentially the same as the pseudo-inverse weight, which is known as an extension of the Hebbian weight (Kanter & Sompolinsky, 1987).
Local targets of DTP. DTP is an improved method of TP, where $\hat{h}_{L}$ is propagated as follows:
$\hat{h}_{i}=g_{l+1}^{\text{diff }}\left(\hat{h}_{l+1},h_{l+1},h_{l}\right)=g_{l+1}\left(\hat{h}_{l+1}\right)-\left(g_{l+1}\left(h_{l+1}\right)-h_{l}\right).$ (S.121)
@@
$\hat{h}_{l}$ $=g_{l+1}\left(\hat{h}_{l+1}\right)-\left(g_{l+1}\left(h_{l+1}\right)-h_{l}\right)$ (S.123)
@@
$=h_{l}-\beta\prod_{i=l+1}^{L}Q_{i}\delta_{L}.$ (S.127)
Therefore, at the equilibrium point for $Q_{l}$, for $l\leq L-2$, we have
$\hat{h}_{l}-h_{l}$ $=-\beta\prod_{i=l+1}^{L}Q_{i}^{*}\delta_{L}$ (S.128)
@@
where $\delta_{L}=\partial\mathcal{L}/\partial f$ and for $l=L-1$, we have $\hat{h}_{L-1}-h_{L-1}=-\beta h_{L-1}(K^{A}_{L}+\mu_{L}^{\prime}I)^{-1}h_{L}^{\top}\delta_{L}$. To avoid an uninteresting change of order, we introduce $\mu_{l}=\mu_{l}^{\prime}/M^{\bar{\mu}_{l}}$ and require that it have the same order as $K_{l}^{A}$. This is essentially equivalent to the valid condition argued in Ishikawa & Karakida (2024), which requires the damping term to have the same order as the preconditioner in the second-order optimization. We note that we can take the ridge-less limit $\mu_{l}^{\prime}\rightarrow 0+$ because $K^{A}_{l}$ (S.39) is typically set to be regular at random initialization in the neural tangent kernel literature (Jacot et al., 2018; Yang, 2020). For instance, this holds true for normalized input samples with $\|x\|=1$.
@@
On Condition A.1. The update for the last layer is identical to that of SGD with BP, thus
$\partial_{\eta^{\prime}}\Delta W_{L}h_{L-1,1}\big{|}_{\eta^{\prime}=0}$ $=-\frac{1}{M^{\theta_{L}-1}}\beta\delta_{L}K^{A}_{L-1}$ (S.134)
Next, we consider the $L-1$ layer.
@@
$=-\frac{1}{M^{\theta_{L-1}-1}}\beta h_{L-1}(K_{L}^{A}+\mu_{L}^{\prime}I)^{-1}(K_{L}^{A}-M^{-1}h_{L}^{\top}y)K^{A}_{L-1}.$ (S.137)
@@
$=-\frac{1}{M^{\theta_{l}-1}}\beta h_{l}\prod_{i=l+1}^{L-1}(h_{i}^{\top}h_{i}+\mu_{i}I)^{-1}h_{i}^{\top}h_{i}(h_{L}^{\top}h_{L}+\mu_{L}I)^{-1}h_{L}^{\top}\delta_{L}K^{A}_{l-1}$ (S.139)
$=-\frac{1}{M^{\theta_{l}-1}}\beta h_{l}\prod_{i=l+1}^{L-1}(K^{A}_{i}+\mu_{i}^{\prime}I)^{-1}K^{A}_{i}(K^{A}_{L}+\mu_{L}^{\prime}I)^{-1}(K^{A}_{L}-M^{-1}h_{L}^{\top}y)K^{A}_{l-1}$ (S.140)
@@
$\partial_{\eta^{\prime}}(W_{L,0}\Delta h_{L-1,1})\big{|}_{\eta^{\prime}=0}$ $=\frac{1}{M^{\theta_{L-1}}}W_{L,0}\text{diag}(\phi^{\prime}_{L-1})(\hat{h}_{L-1}-h_{L-1})h_{L-2}^{\top}$ (S.144)
$=-\frac{1}{M^{\theta_{L-1}-1}}\beta h_{L}(h_{L}^{\top}h_{L}+\mu_{L}^{\prime}I)^{-1}h_{L}^{\top}\delta_{L}K^{A}_{L-2}$ (S.145)
$=-\frac{1}{M^{\theta_{L-1}-1}}\beta h_{L}(K^{A}_{L}+\mu_{L}^{\prime}I)^{-1}(K^{A}_{L}-M^{-1}h_{L}^{\top}y)K^{A}_{L-2}.$ (S.146)
@@
C.1.1 Disappearance of Kernel regime
Corollary C.1.
@@ C.1.1 Disappearance of Kernel regime
Note that, precisely speaking, $r_{L-1}=0$ does not necessarily imply $r_{l<L-1}=0$. However, it is often considered unnatural (or uninteresting) to examine cases in which the progress of learning depends on individual layers. Therefore, the $\mu$P typically assumes a uniform parameterization, meaning $r_{l<L}=r$ (see Theorem G.4 of Yang & Hu (2021)). In this sense, $r_{L-1}=0$ indicates the disappearance of the kernel regime.
One might be surprised by the fact that $b_{L}=1/2$ is allowed in the feature learning and that the kernel regime disappears. Note that the feedback weight in the last layer (S.120) essentially differs from $W_{L}$. The feedback weight receives $h_{L}$ as input whereas $W_{l}$ receives $h_{L-1}$. This makes
@@
$W_{L}\sim 1/M^{a_{L}+b_{L}}$. The gradient is proportional to $Q_{L}$ in TP and $W_{L}$ in BP. The feedback weight contributes more significantly to TP’s gradient when $a_{L}+b_{L}\geq 1/2$. This eventually makes the index $r$ of the hidden layer (S.143) get quite large even for $a_{L}+b_{L}=1/2$. We also need to be careful about the order of condition A.2 (S.148). Because the feedback weight (S.120) has a lower alignment exponent (Everett et al., 2024), this causes the condition 2 of TP to be smaller than that of SGD (or K-FAC), i.e., $1/M^{r_{L-1}+(a_{L}+b_{L})-1}$. Therefore, stable feature learning is possible even for $a_{L}+b_{L}=1/2$.
@@
Substituting these into Eqs. (S.137,S.140), we obtain
@@
C.2 Stable parameterization for feedback network
@@
where dividing by $M_{L-1}$ is to ensure that $L(Q_{l})=\Theta(1)$, which is the default setting in PyTorch. We consider the parameterization in the feedback network:
@@
Appendix D Additional Experiments
D.1 Predictive Coding
D.1.1 Linear Network
Figure S.1 : During the training of the linear network, it converges to the analytical solution. We trained a 3-layer linear network using synthetic data.